### v1

# The simplest Kochen-Specker set

Abstract
Kochen-Specker (KS) sets are fundamental in physics. Every time nature produces bipartite correlations attaining the nonsignaling limit, or two parties always win a nonlocal game impossible to always win classically, is because the parties are measuring a KS set. The simplest quantum system in which all these phenomena occur is a pair of three-level systems. However, the simplest KS sets in dimension three known are asymmetrical and require a large number of bases (the current minimum is 16, set by Peres and Penrose). Here we present a KS set that is much more symmetrical and easier to prove than any previous example. It sets a new record for minimum number of bases, 14, and enables us to refute Conjecture 2 in Phys. Rev. Lett. 134, 010201 (2025), setting a new record for qutrit-qutrit perfect strategies with a minimum number of inputs: 5-9.
Introduction.—A Kochen-Specker (KS) set [1] is a finite set of unit vectors $\mathcal{V}$ (or, equivalently, observables represented by rank-one projectors) in a Hilbert space ${\cal H}=\mathbb{C}^{d}$ of finite dimension $d\geq 3$, which does not admit an assignment $f:\mathcal{V}\rightarrow\{0,1\}$ such that $f(u)+f(v)\leq 1$ for $u,v\in\mathcal{V}$ orthogonal, and $\sum_{u\in b}f(u)=1$ for every orthonormal basis $b\subseteq\mathcal{V}$.
KS sets are fundamental in physics for many reasons. They prove the impossibility of noncontextual hidden-variable theories [1]. They are at the basis of the experimental tests of nature’s state-independent contextuality [2, 3, 4, 5, 6], and of experiments producing nonlocality and contextuality simultaneously [7, 8]. KS sets allow contextuality to be converted into nonlocality [9]. KS sets unique up to unitaries can be certified from their input-output statistics using any state of full rank [10] and enable to Bell self-test supersinglets of any dimension [11]. KS sets are necessary [12] for bipartite correlations attaining the nonsignaling limit [13, 14, 15, 16], and, via a result in [17], for bipartite perfect strategies (aka quantum pseudotelepathy) [18], and bipartite correlations with maximal nonlocal fraction (aka fully nonlocal correlations) [19]. In these last two roles, KS sets are crucial ingredients in some important results in quantum foundations and quantum computation, including the resolution of Tsirelson’s problem [20] and the quantum computational advantage in shallow circuits [21].
There are beautiful small KS sets in dimensions four [22] and higher [23]: they are symmetrical, the impossibility of a KS assignment is easy to prove, and each vector has nonzero components of equal magnitude. In contrast, none of the known small KS sets of dimension three, i.e., Schütte-33 [24] (the number indicates the number of vectors), Peres-33 [25], Conway-Kochen-31 [26][p. 114], and Penrose-33 [27], has any of these properties.
[…]
Minimal-basis KS set.—Our set has 33 vectors. Their components and orthogonality relations are shown in Fig. 1. The proof that they do not admit a KS assignment is as follows. The 13 vectors in the central triangle in Fig. 1 are the Yu-Oh set [28] (see also [29][Fig. 2.9]), which admits KS assignments only if at most one of the four non-white nodes is assigned $1$. The fact that no KS assignment is possible for all vectors in Fig. 1 follows from the observation that any assignment would require exactly one of the red (or green, or blue, or yellow) nodes to be assigned $1$ because the three red (blue, green, yellow) nodes are mutually orthogonal. However, that is impossible because there is only room for three $1$’s (one per Yu-Oh set), while there are four colors (red, green, blue, and yellow).
As frequently pointed out (see, e.g., [23] and references therein), the true measure of physical simplicity of a KS set is its number of bases. This has been corroborated by the recent discovery that, for the simplest KS sets, what determines the sum of Alice and Bob’s inputs necessary to attain a perfect quantum strategy is the number of complete bases [12, 30]. The original KS set [1] had 43 complete bases [31][Fig. 1]. Peres set the current record, 16 complete bases, in 1991 [25]. Penrose equaled the record in 2000 with a different construction [27]. Conway-Kochen-31 [26][p. 114] has 17 complete bases and Schütte-33 [24] has 20 complete bases. The set introduced in this Letter sets a new record: 14. They correspond to the large triangle in Fig. 1 (which is repeated three times), the 9 small triangles, and the four colored bases.
Importantly, the KS introduced here refutes Conjecture 2 in [12], which stated that the qutrit-qutrit perfect strategy with minimum $|X|\times|Y|$, where $|X|$ is Alice’s input cardinality and $|Y|$ is Bob’s, has $|X|=7$ and $|Y|=9$, and corresponds to distributing the 16 bases of Peres-33 (or Penrose-33) in a particular way. Now we can formulate the following.
Conjecture: The qutrit-qutrit perfect strategy with minimum $|X|\times|Y|$ has $|X|=5$ and $|Y|=9$, and is produced by preparing a maximally entangled qutrit-qutrit state, and allowing Bob to measure the nine bases corresponding to the nine small triangles in Fig. 1, and Alice to measure the other five complete bases in Fig. 1.
That, given the set in Fig. 1, the protocol described above produces a perfect quantum strategy of minimal $|X|\times|Y|$ follows from applying the method in [30]. Remarkably, the symmetries of the KS set dictate the precise distribution of bases required to achieve minimality.
Conclusions.—For almost 35 years, Peres-33 [25] has been the “standard” KS set: the one with the fewest bases, the one with the highest degree of symmetry, the one that determines the simplest scenario in which nature reaches the nonsignaling limit, produces perfect nonlocal strategies, and generates fully nonlocal correlations using a bipartite system of minimal dimension. Peres-33 has been the default option for a KS set. It is the one that is in the cover of modern books [32]. It is the one used to explain the use of KS sets in nonlocal games [33]. To the point that even Conway and Kochen used Peres-33 (instead of Conway-Kochen-31) to prove both the free-will theorem [34] and the strong free-will theorem [35].
In this Letter, we have introduce a KS set that substantially improves Peres-33 in all aspects. Table 1 enables us to appreciate the improvement. In a nutshell, our result shows that nature does it better than we thought: A KS state-independent contradiction with noncontextual hidden-variable theories only requires 14 bases. They define a much more elegant (symmetrical) KS set than any previous one in dimension three. Nature only needs Alice to choose between $5$ settings and Bob to choose between $9$ settings to produce, with pairs of systems of minimal dimension, correlations attaining the nonsignaling limit, allowing them to always win a game impossible to always win with classical methods, and showing full nonlocality (i.e., nonlocal fraction $1$).
Number of bases Degree of symmetry Minimal bipartite perfect strategy
Schütte-33 [24] 2020 88 898-9
Conway-Kochen-31 [26] $17$ 44 898-9
Peres-33 [25] $16$ $48$ 797-9
Penrose-33 [27] $16$ $48$ 797-9
This Letter $14$ $144$ 595-9
Table 1: Simplest KS sets in dimension three. The degree of symmetry is quantified by the size of the automorphism group (i.e., all permutations of the vertices that preserve adjacency) of the orthogonality graph. The minimal bipartite perfect strategy is the perfect strategy that minimizes $|X|\times|Y|$, where |X||Y||X|-|Y| indicates Alice’s and Bob’s number of settings, respectively. The minimal perfect strategies have been obtained using the method in [30].
The fact that it has taken us almost 35 years to make progress and that the remaining margin for improvement is shrinking thanks to new methods for exploring bipartite perfect quantum strategies [17, 12, 30] and finding lower bounds on the size of the smallest KS set [36, 37] may indicate that we are already at (or very close to) the true natural limit.
I thank Stefan Trandafir and Zhengyu Li for helpful discussions. I also thank Stefan Trandafir for computing the sizes of the automorphism groups. This work was supported by EU-funded project FoQaCiA.